OpenStudy (anonymous):

if you make a composite function g(g(x)) where g(x) = x^3, would the answer be x or x^9? and explain your reasoning please...

OpenStudy (anonymous):

it would be \(g(g(x))=g(x^3)=(x^3)^3=x^9\) what you said second

OpenStudy (anonymous):

i also thought that but my text book says its x for some reason and i dont understand why

hero (hero):

lol

OpenStudy (anonymous):

?

hero (hero):

Sorry. Your book is wrong

OpenStudy (anonymous):

okay agreed :)